The Centers for Disease Control reported the percentage of people 18 years of age and older who smoke (CDC website, DecemberPlease show step by step on EXCEL
Answer #1

(a) Here as it is given that margin of error = 0.02

with 95% confidence.

so here first we will find critical test statistic for 95% confidence i.e. = NORMSINV(0.5 + 0.95/2) = NORMSINV(0.975) = 1.96 (Using Excel)

so here

margin of error = critical test statistic * standard error of sample proportion

standard error of sample proportion = sqrt [p * (1-p)/n]

here given p = 0.26

standard error = sqrt [0.26 * (1-0.26)/n]

0.02 = 1.96 * sqrt [0.26 * (1 - 0.26)/n]

n = 1847.81 or 1848

(b) Here as given point estimate = p^ = x/n = 520/1848 = 0.2814

(c) Now we would require to calculate 95% confidence interval

95% confidence interval = p^ +- MOE

Margin of error = Critical Test statistic * standard error of proportion

Critical Test statistic for 95% confidence interval = NORMSINV(0.5 + 0.95/2) = NORMSINV(0.975) = 1.96

Standard error of proportion = sqrt [0.2814 * (1 - 0.2814/1848] = 0.0105

Margin of error = 1.96 * 0.0105 = 0.0205

95% confidence interval = 0.2814 +- 0.0205 = (0.2609, 0.3019)